OpenStudy (idealist10):

Find an integrating factor of the form u(x, y)=P(x)Q(y) for (3x^2*y^3-y^2+y)dx+(-xy+2x)dy=0.

OpenStudy (idealist10):

@SithsAndGiggles

ganeshie8 (ganeshie8):

*

OpenStudy (anonymous):

u(x,y)M dx + u(x,y)N dy=0 Now, the above equation is an exact differential equation. \[\frac{ \delta }{ \delta y }(u(x,y) M)=\frac{ \delta }{ \delta x }(u(x,y) N)\]

OpenStudy (anonymous):

Use this equation to find u(x,y).

ganeshie8 (ganeshie8):

looks the solution is going to be very long

OpenStudy (anonymous):

Indeed.

ganeshie8 (ganeshie8):

it will be easy if any one of the functions P or Q is 1

ganeshie8 (ganeshie8):

i worked these problems assuming the IF is a function of x or y alone but never worked a problem that requires IF to be function of both variables

OpenStudy (ikram002p):

hmm partial differential equation hehe it hates me lol xD

ganeshie8 (ganeshie8):

its ODE only... we are finding IF to make it exact

OpenStudy (ikram002p):

u(x,y ) partial

ganeshie8 (ganeshie8):

yeah requires taking partial derivatives, but this is not a paritial differential eqn

OpenStudy (ikram002p):

sure ?

ganeshie8 (ganeshie8):

ugh u don't see any partial derivatives in the given eqn

ganeshie8 (ganeshie8):

*i

OpenStudy (ikram002p):

im not that good , but i put this question on my list to learn , tomorrow i'll type what ever i got xD

OpenStudy (anonymous):

Using the method we've discussed before: \[\frac{u_x}{u}N-\frac{u_y}{u}M=M_y-N_x\] We want to find \(F(x)=\dfrac{u_x}{u}\) and \(G(y)=\dfrac{u_y}{u}\). Partials: \[M_y=\frac{\partial }{\partial y}[3x^2y^3-y^2+y]=9x^2y^2-2y+1\\ N_x=\frac{\partial }{\partial x}[-xy+2x]=-y+2\] Substituting: \[\begin{align*}F(x)(-xy+2x)\\-G(y)(3x^2y^3-y^2+y)&=9x^2y^2-2y+1-(-y+2)\\\\ G(y)&=\frac{F(x)(-xy+2x)-9x^2y^2+y+1}{3x^2y^3-y^2+y}\end{align*}\] Let's try to eliminate the factor of \(3x^2y^3-y^2+y\) by finding a function \(C(y)\) such that \[F(x)(-xy+2x)-9x^2y^2+y+1=C(y)(3x^2y^3-y^2+y)\] Supposing \(C(y)=-\dfrac{3}{y}\), we have \[\begin{align*}F(x)(-xy+2x)-9x^2y^2+y+1&=-9x^2y^2+3y-3\\\\ F(x)(-xy+2x)&=2y-4\\\\ F(x)&=\frac{2y-4}{-x(y-2)}\\\\ F(x)&=-\frac{2}{x}\end{align*}\] Solving for \(G(y)\): \[\begin{align*}G(y)&=\frac{-\dfrac{2}{x}(-xy+2x)-9x^2y^2+y+1}{3x^2y^3-y^2+y}\\\\ G(y)&=\frac{-9x^2y^2+3y-3}{y(3x^2y^2-y+1)}\\\\ G(y)&=\frac{-3(3x^2y^2-y+1)}{y(3x^2y^2-y+1)}\\\\ G(y)&=-\frac{3}{y} \end{align*}\] So the IF will be \[\begin{align*}\log u(x,y)&=\int F(x)~dx+\int G(y)~dy\\\\ &=-2\int\dfrac{dx}{x}-3\int\frac{dy}{y}\\\\ u(x,y)&=\frac{1}{x^2y^3}\end{align*}\]

OpenStudy (anonymous):

@Idealist10 there's a slight deviation from the approach I described the first time around. Before we let \(C\) be a constant, while here \(C\) is a function (though even when it is a constant, it's still technically a function). The point is that this sort of "greedy" approach lets you simplify the PDE so that you can find those one-dimensional functions \(F\) and \(G\).

OpenStudy (idealist10):

Okay, I get it. Thanks a lot!